OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: x + y = 6 Line B: x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There is no solution. It is (6, 4). It is (4, 6).

OpenStudy (carolinar7):

Start by eliminstion which I think will be easier

OpenStudy (anonymous):

I think its no solution

OpenStudy (carolinar7):

*elimination

OpenStudy (carolinar7):

Yes you are right they are parrallel lines

OpenStudy (anonymous):

The distances (y), in miles, of two cars from their starting points at certain times (x), in hours, are shown by the equations below: Car A y = 55x + 32 Car B y = 42x + 58 After how many hours will the two cars be at the same distance from their starting point and what will that distance be? 2 hours, 142 miles 2 hours, 145 miles 3 hours, 142 miles 3 hours, 145 miles

Vocaloid (vocaloid):

since both equations are in the form y = ____, we can set them equal to each other 55x + 32 = 42x + 58 solve for x

OpenStudy (anonymous):

I think it's 2,142

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

Variable p is 2 more than variable d. Variable p is also 1 less than variable d. Which pair of equations best models the relationship between p and d? p = d + 2 p = d − 1 p = d − 2 p = d + 1 d = 2p d = 2p − 1 d = 2p d = 2p + 1

OpenStudy (anonymous):

I think its the first one

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

A set of equations is given below: Equation A: e = 4f + 1 Equation B: e = 3f + 5 Which of the following is a step that can be used to find the solution to the set of equations? 4f + 1= 3f 4f = 3f + 5 4f + 5 = 3f + 1 4f + 1 = 3f + 5 I hope your right because this is a final exam

OpenStudy (triciaal):

we cannot help on an exam it is considered cheating

OpenStudy (confluxepic):

"Posting questions from tests, exams or other third party content may be an infringement of copyrights of the owner of the posted content. In most cases, asking for help on an exam or test, or helping someone during their test or exam, will be considered cheating by the institution or by OpenStudy. OpenStudy reserves the right to remove such content and/or suspend, ban or otherwise remove Users from the system who are suspected of cheating. If you feel there is an error, You may email contact@openstudy with an explanation but You acknowledge that the determination to remove content or suspend, ban or otherwise remove a User from the system is within the sole discretion of Company." http://openstudy.com/terms-and-conditions

OpenStudy (confluxepic):

"By using the OpenStudy Site or Service, you acknowledge that OpenStudy cooperates fully with schools, colleges and other educational institutions (together “Schools”) and reserves the right to disclose User information to a School if there is an allegation by that School of cheating or violation of the School’s academic policies."